# Leopard[Uncaptioned image]: A Vision Language Model for Text-Rich Multi-Image Tasks

### v2

@@ 3.2 Adaptive High-resolution Multi-image
To overcome these limitations, a natural solution is dividing a high-resolution image into multiple smaller sub-images, each of which is independently processed by the model’s visual encoder [40, 13]. This partitioning allows for the extraction of more fine-grained visual details, making it possible to capture small or densely packed textual elements. However, a major drawback of this approach is that it significantly increases the length of visual feature sequence. When applied to scenarios involving multiple image inputs, the feature sequences are easily exceeding the model’s maximum sequence length limit. To address the issue, we follow the image-splitting idea and propose a novel adaptive high-resolution multi-image encoding strategy as follows.
[…]
@@
Image Partitioning: For each image, we perform a grid search over possible number of rows $r$ and columns $c$ (where $1\leq r,c\leq S^{\prime}_{i}$ and $r\times c\leq S^{\prime}_{i}$) to find the optimal cropping configuration that maximizes the effective resolution within the allocated sub-images [32]. This configuration results in the original image being padded and resized to a target resolution of $(h^{\prime}_{i}=r\times v,w^{\prime}_{i}=c\times v)$. We then divide the image into $r\times c$ sub-images of size $(v\times v)$. Additionally, the original image is directly resized to $(v\times v)$, which provides a global view of the visual content.
@@ 4 Experiment
Models Text-Rich Multi-Image Text-Rich Single-Image
MVQAD DUDE SlideVQA MCQA MH Multi Avg. VQAT VQAD VWB Avg.
Otter-9B 0.17 0.15 5.95 1.08 0.14 1.50 23.18 3.53 10.20 12.30
Emu2-Chat 17.58 13.79 0.60 2.40 0.72 7.02 66.60 5.44 18.17 30.07
MM1-7B-Chat - - - - - - 72.80 - - -
VILA-LLaMA3-8B 30.75 19.75 24.72 1.87 3.66 16.15 66.30 30.38 23.37 40.02
mPlug-DocOwl-1.5 35.85 16.94 4.54 0.26 0.86 11.69 68.60 82.20 29.80 60.20
Idefics2-8B 46.67 23.06 25.14 2.59 9.89 21.47 70.40 67.30 23.76 53.82
LLaVA-NeXT-Inter 39.92 24.04 23.46 14.34 3.55 21.06 62.76 75.70 21.36 53.27
Mantis-LLaVA 31.89 17.73 16.81 9.72 3.46 15.92 59.20 39.02 17.88 38.70
Mantis-Idefics2 51.61 27.74 24.02 12.97 5.48 24.36 63.50 54.03 22.47 46.67
Leopard-LLaVA 53.90 35.94 23.83 9.68 10.76 26.82 67.70 68.07 24.91 53.56
Leopard-Idefics2 66.06 40.74 34.93 18.03 10.09 33.97 80.40 74.79 25.60 60.26
Table 3: Experiment results of baseline models and Leopard on 8 benchmarks of text-rich images. We use abbreviated benchmark names due to space limits. MVQAD: Multi-page DocVQA, MCQA: MultiChartQA, MH: MultiHiertt, VQAT: TextVQA, VQAD: DocVQA, VWB: VisualWebBench. Following [64], for MVQAD, DUDE, and VQAD, we use average normalized levenshtein similarity [5] as the metric. For other benchmarks, accuracy is used as the metric, which measures if the predicted answer exactly matches any target answer.
@@ 4.1 Implementation Details
Model Architecture. We train our models on two base architectures: LLaVA [39] and Idefics2 [28]. For Leopard-LLaVA, we use SigLIP-SO-400M [73] with $364\times 364$ image resolutions as the visual encoder since it supports larger resolution than the commonly used $224\times 224$ resolution CLIP visual encoder [55]. Each image is encoded into a sequence of $26\times 26=676$ visual features under a patch size of $14$. With the visual feature pixel shuffling strategy, each image is further processed into a sequence of $169$ visual features. We limit the maximum number of images ($M$) in each sample to $50$, which produces up to $8,450$ visual features in total. Following llava [39], we adopt a two-layer MLPs as the visual-language connector. We use LLaMA-3.1 [53] as the LM.
For Leopard-Idefics2, we follow the architecture of Idefics2-8B which uses SigLIP-SO-400M as the visual encoder but increases its image resolution to $980\times 980$ to make the text legible. The features outputted by the visual encoder are compressed with a feature resampler into $64$ tokens per image. Idefics2-8B adopts the Mistral-7B [20] as the LM.
Training Details. When training Leopard-LLaVA, we first train the visual-language connector using LLaVA’s 558K multimodal pre-training dataset. Subsequently, we fine-tune the model (with both the connector and the LM unfrozen) using our Leopard-instruct data. As for Leopard-Idefics2, it is pre-trained on a dataset comprised of over 350M multimodal samples. Given the computational challenges of reproducing such extensive pre-training, and to ensure a fair comparison with baselines that utilize the pre-trained Idefics2 checkpoint, we directly adopt Idefics2’ visual feature resampler and fine-tune the model on the Leopard-instruct dataset.
We train both Leopard-LLaVA and Leopard-Idefics2 on 64 A100-40G GPUs with a global batch size of 128. We use the AdamW optimizer with $\beta_{1}=0.9$, $\beta_{2}=0.999$. Following [21], we use a learning rate of $1\times 10^{-5}$ for Leopard-LLaVA and $5\times 10^{-6}$ for Leopard-Idefics2 to protect its pretrian knowledge. We use a cosine learning rate scheduler with a linear learning rate warm-up for the first $3\%$ steps. All model variants are trained 1 epoch under the same hyperparameters. It takes around 120 GPU days to train Leopard under both settings.
@@ 4.2 Baseline Models
We compare Leopard against a range of existing open-source MLLMs that support multi-image inputs. These baseline models include Otter-9B [30], Emu2-Chat-34B [61], MM1-7B-Chat [52], Mantis [21], VILA [38], Idefics2-8B [28], and LLaVA-NeXT-Interleave [33].
Models MIRB MiBench MMMU MathVista SQAI Avg.
Otter-9B 20.74 43.72 30.89 22.00 60.44 35.55
Emu2-Chat 36.02 58.93 34.10 30.40 65.69 45.03
MM1-7B-Chat - - 37.00 35.90 72.60 -
VILA-LLaMA3-8B 40.87 53.70 36.90 35.40 79.90 49.35
mPlug-DocOwl-1.5 25.39 40.80 35.44 29.50 64.40 39.11
Idefics2-8B 33.02 46.39 42.90 45.00 89.04 51.27
LLaVA-NeXT-Inter 44.38 74.52 38.44 32.10 72.63 52.41
Mantis-LLaVA 40.76 59.96 40.10 34.40 74.90 50.02
Mantis-Idefics2 41.80 56.80 41.10 40.40 81.30 52.28
Leopard-LLaVA 42.00 60.80 43.00 45.50 85.57 55.37
Leopard-Idefics2 41.38 61.74 40.11 44.80 90.38 55.68
Table 4: Experimental results on general domain benchmarks. We abbreviate the $\mathrm{Image}$ split of ScienceQA as SQAI.
Models that only support a single image input are excluded from our comparisons, except for mPlug-DocOwl-1.5 [16], as it is primarily trained on visual document data and demonstrates strong capabilities on text-rich image tasks. Table 2 demonstrates a detailed comparison of the model training details of between baseline models and our proposed Leopard, which highlights their architecture, image resolution and training data differences.
@@ 4.3 Evaluating Benchmarks
We evaluated Leopard and baseline methods across three categories of vision-language tasks on (1) single text-rich image evaluation, (2) multiple text-rich images evaluation, and (3) general reasoning evaluation. Benchmarks for (1) include TextVQA [59], DocVQA [51], and VisualWebBench [43]. Benchmarks for (2) include Multi-page DocVQA [64], DUDE [25], SlideVQA [62], Multihiertt [77], and MultiChartQA [2], which cover a diverse range of typical multi-image tasks, such as document understanding and slide question answering. Benchmarks for (3) include MMMU [71], MathVista [45], ScienceQA [56], MIRB [76] and MiBench [42], which evaluate MLLMs from different perspectives, including world knowledge, mathematics, and scientific reasoning etc.
Ablation Settings Text-Rich Multi-Image Text-Rich Single General
MVQAD DUDE SlidesVQA Multi Avg. TextVQA DocVQA MMMU MathVista
($\star$) Our Best Setting (as in Table 3): LLaMA-3.1 + Adaptive + 1 1 1
Leopard-LLaVA 53.90 35.94 23.83 37.89 67.70 68.07 43.00 45.50
(1) Effect of Adaptive High-Resolution Encoding: LLaMA-3.1 + 1 1 1
- w/o Adaptive 40.44 26.16 20.93 29.17(8.7$\downarrow$) 60.18 44.69 41.00 42.40
(2) Effect of Backbone LLMs: LLaMA-3 + Adaptive + 1 1 1
- with LLaMA-3 .1 48.66 32.64 25.75 35.68(2.2$\downarrow$) 67.08 54.92 41.22 42.10
(3) Effect of Data Domains: LLaMA-3.1 + Adaptive
- with chart web 43.79 29.50 23.10 32.13(5.7$\downarrow$) 66.78 56.60 40.67 44.80
- with doc web 54.33 35.65 18.73 36.23(1.7$\downarrow$) 66.86 50.78 41.89 39.60
- with doc chart 54.62 35.70 20.79 37.02(0.9$\downarrow$) 67.40 67.82 41.78 44.00
Table 5: Ablation studies on Leopard-LLaVA from four different perspectives: (1) evaluating the impact of Adaptive High-Resolution Encoding, (2) pre-training LLaVA by initializing with checkpoints from either LLaMA-3 or LLaMA-3.1, and (3) examining the impact of using different data domains for instruction tuning, including doc, chart, and web.
@@ 4.4 Main Experimental Results
Leopard achieves outstanding performance on text-rich, multi-image benchmarks, as shown in Table 3. Notably, both Leopard-LLaVA and Leopard-Idefics2 significantly outperform all baselines. Leopard-Idefics2 becomes the strongest open-source MLLM in this area, achieving an average improvement of 9.61 points over the previous best performance.
In single-image text-rich scenarios, Leopard outperforms several recent strong models, including VILA and LLaVA-NeXT. Leopard even achieves slightly higher average scores than the state-of-the-art mPlug model, despite mPlug being trained on 4M single-image data while Leopard is tuned on <200K. This demonstrates that training on multi-image data from Leopard-instruct also benefits model performance on single-image tasks.
@@ 4.5 More Analysis
Leopard-instruct mainly cover three main domains, i.e. documents & slides ( doc), tables & charts ( chart), and websites ( web). To assess the impact of data from different domains, we conduct ablation studies on three variants of Leopard, with the results presented in Table 5 Removing any part of the training data results in performance degradation. The most significant drop occurs when we exclude document data while removing web data leads to a slight decrease. However, the mixed-domain datasets, such as LLaVAR and mPlugDocReason, also contain data in these domains which are challenging to isolate and ablate. This may contribute to the relatively preserved performance even after the ablation of certain data sources.
@@ 5 Conclusion
In this paper, we introduce Leopard, a novel MLLM specifically designed for text-rich, multi-image tasks. Leopard is equipped with two key innovations: (1) Leopard-instruct, a large-scale instruction-tuning dataset that encompasses a wide range of text-rich, multi-image instructions, and (2) an adaptive image encoding module capable of processing multiple high-resolution images efficiently. Our experimental results across diverse benchmarks highlight Leopard’s superior performance compared to existing open-source MLLMs, particularly in text-rich multi-image scenarios. Further analysis and ablation studies underscore the effectiveness of both the collected dataset and adaptive encoding strategy, solidifying Leopard’s contribution to multimodal research.
@@ A.1 Leopard-Instruct
To train Leopard, we created a large instruction-tuning dataset, Leopard-instruct, with 925K instances, including 739K designed for text-rich, multi-image scenarios. Despite surveying existing datasets, we found only 154K suitable text-rich, multi-image samples – insufficient for effective instruction tuning, which is far from sufficient for effective instruction tuning, as shown in prior MLLM studies [21, 28, 33]. To overcome this limitation, we developed several data collection pipelines to collect high-quality text-rich, multi-image data, resulting in additional 585K instances.
@@
ArxivQA [34] Doc Reformed 1-3 Existing 81
DUDE [25] Doc Public 1-50 Augmented 23
MP-DocVQA [64] Doc Public 1-20 Augmented 36
DocVQA [51] Doc No 1 None 39
TAT-DQA [79] Doc Reformed 2-5 Augmented 13
SlidesGeneration [57] Slides Repurposed 1-20 Augmented 3
SlidesVQA [62] Slides Public 20 Augmented 10
@@
Multihiertt [77] Table Public 3-7 Existing/Augmented 15
MultiTabQA [54] Table Public 1-2 Augmented 6
TableGPT [35] Table Split 2 Existing 4
TabMWP [46] Table No 1 Existing 23
ChartGemma [49] Chart Reformed 1-4 Existing 65
DVQA [22] Chart Reformed 1-3 None 200
FigureQA [23] Chart Reformed 1-2 None 36
ChartQA [48] Chart Reformed 2 Augmented 32
@@
Mind2Web [12] Web Split 1-5 None 7
WebsiteScreenshots [4] Web No 1 Augmented 2
Omniact [24] Web No 1 None 1
RICO [15] Web Reformed 1-4 None 25
WebVision [36] Web No 1 Existing 1
WebUI [68] Web No 1 None 19
LLaVAR [75] Mix No 1 Existing 15
MathV360k [58] Mix No 1 None 38
Monkey [37] Mix Reformed 1-3 None 92
MPlugDocReason [16] Mix No 1 Existing 25
IconQA [47] Other Public 1-6 Augmented 64
InfographicVQA [50] Other No 1 Augmented 23
MapQA [6] Other Reformed 1-2 None 4
@@
A.2 Prompts

### v3

@@ 3.2 Adaptive Resolution Multi-Image Encoding
To address these limitations, a viable solution is to divide a high-resolution image into smaller sub-images, each processed independently by the model’s visual encoder (Liu et al., 2024a; Dong et al., 2024). This partitioning captures finer visual details, enabling recognition of small or densely packed text. However, this approach greatly extends the visual feature sequence length, often surpassing the model’s maximum sequence limit when handling multiple images. To mitigate this, we introduce an adaptive high-resolution multi-image encoding strategy.
Image Allocation Computing: To prevent the number of sub-image visual features from exceeding the LLM’s maximum sequence length, we first set a budget $M$ for the total number of sub-images. We allocate this budget proportionally to each input image based on their original sizes. For each image $\mathbf{i}$ with dimensions $h_{i}\times w_{i}$, we calculate the initial number of sub-images $S_{i}$ as:
@@
Image Partitioning: For each image, we perform a grid search over possible number of rows $r$ and columns $c$ (where $1\leq r,c\leq S^{\prime}_{i}$ and $r\times c\leq S^{\prime}_{i}$) to find the optimal cropping configuration that maximizes the effective resolution within the allocated sub-images (Li et al., 2024a). This configuration results in the original image being padded and resized to a target resolution of $(h^{\prime}_{i}=r\times v,w^{\prime}_{i}=c\times v)$. We then divide the image into $r\times c$ sub-images of size $(v\times v)$. Additionally, the original image is directly resized to $(v\times v)$, which provides a global view of the visual content.
Models Text-Rich Multi-Image
MVQAD DUDE SlideVQA MCQA MH Average
Otter-9B 0.17 0.15 5.95 1.08 0.14 1.50
Emu2-Chat 17.58 13.79 0.60 2.40 0.72 7.02
VILA-LLaMA3-8B 30.75 19.75 24.72 1.87 3.66 16.15
mPlug-DocOwl-1.5 35.85 16.94 4.54 0.26 0.86 11.69
Idefics2-8B 46.67 23.06 25.14 2.59 9.89 21.47
LLaVA-NeXT-Inter 39.92 24.04 23.46 14.34 3.55 21.06
Llama-3.2-11B 57.60 20.77 19.43 6.25 3.16 21.44
Qwen2-VL-7B 71.86 40.86 19.08 6.89 7.37 29.21
Mantis-LLaVA 31.89 17.73 16.81 9.72 3.46 15.92
Mantis-Idefics2 51.61 27.74 24.02 12.97 5.48 24.36
Leopard-LLaVA 57.56 37.30 27.53 14.30 7.70 28.88
Leopard-Idefics2 66.06 40.74 34.93 18.03 10.09 33.97
Leopard-LLaVA-Pro 70.60 43.82 37.51 24.30 12.12 37.67
     – compared to SoTA Qwen2-VL (-1.3) (+3.0) (+18.4) (+17.4) (+4.8) (+8.5)
Table 4: Experiment results of baseline models and Leopard on 8 benchmarks of text-rich images. We use abbreviated benchmark names due to space limits. MVQAD: Multi-page DocVQA, MCQA: MultiChartQA, MH: MultiHiertt. Following (Tito et al., 2022), for MVQAD and DUDE, we use average normalized Levenshtein similarity (ANLS) as the metric. For other benchmarks, accuracy is used as the metric, which measures if the predicted answer exactly matches any target answer. We highlight the best and the second best results with Bold and underline, respectively.
Image Encoding: Most vision encoders transform an image into a sequence of visual features $\mathbf{v}\in\mathbb{R}^{L\times d}$, where $L$ denotes the sequence length and $d$ the feature dimension. Typically, $L$ is in the hundreds; for instance, the SigLIP encoder yields $L=676$ and $d=1152$ for an input image. Given that most LLMs have a sequence length limit of 8K tokens, this restricts the number of encoded images to at most 12 without text input, severely limiting image capacity. To address this, inspired by pixel shuffling (Chen et al., 2024; Laurençon et al., 2024a), we concatenate $n$ adjacent visual features along the feature dimension, reducing $L$ by a factor of $n$. This results in a compressed visual feature sequence $\mathbf{v}^{\prime}\in\mathbb{R}^{\frac{L}{n}\times nd}$, effectively allowing more images to be accommodated.
@@ 4 Experiment
Ablation Settings Backbone Adaptive CoT Text-Rich Multi-Image Text-Rich Single General
MVQAv2 DUDE SlidesVQA Avg. TextVQA DocVQA MMMU MathVista
Leopard-LLaVA LLaMA-3.1 57.56 37.30 27.53 40.80 67.70 68.07 43.00 45.50
– w/o Adaptive LLaMA-3.1 $\times$ 40.44 26.16 20.93 29.17($\downarrow$11.6) 60.18 44.69 41.00 42.40
– w/o CoT LLaMA-3.1 $\times$ 52.24 34.23 22.57 36.34($\downarrow$4.5) 66.34 64.31 41.78 43.40
– w/ LLaMA-3 LLaMA-3 48.66 32.64 25.75 35.68($\downarrow$5.1) 67.08 54.92 41.22 42.10
Table 5: Ablation study on Leopard-LLaVA: impact of removing or varying Adaptive and CoT components, and of using different backbones.
Ablation Settings Data Text-Rich Multi-Image Text-Rich Single General
size MVQAD DUDE SlidesVQA Multi Avg. TextVQA DocVQA MMMU MathVista
Leopard-LLaVA 925K 57.56 37.30 27.53 40.80 67.70 68.07 43.00 45.50
- w/o doc 720K 43.79 29.50 23.10 32.13(8.7$\downarrow$) 66.78 56.60 40.67 44.80
- w/o chart 524K 54.33 35.65 18.73 36.23(4.6$\downarrow$) 66.86 50.78 41.89 39.60
- w/o web 870K 54.62 35.70 20.79 37.02(3.8$\downarrow$) 67.40 67.82 41.78 44.00
- only existing 340K 49.29 31.96 20.46 33.90(6.9$\downarrow$) 60.22 60.58 40.67 40.40
Table 6: Data ablation studies on Leopard-LLaVA: evaluating the impact of different data domains for instruction tuning, including doc, chart, and web, as well as using only existing data listed in Table 1.
@@ 4.1 Implementation Details
Model Architecture. We train our models on two base architectures: LLaVA (Liu et al., 2023a) and Idefics2 (Laurençon et al., 2024b). For Leopard-LLaVA, we use SigLIP-SO-400M (Zhai et al., 2023) with $364\times 364$ image resolutions as the visual encoder since it supports larger resolution than the commonly used $224\times 224$ resolution CLIP visual encoder (Radford et al., 2021). Each image is encoded into a sequence of $26\times 26=676$ visual features under a patch size of $14$. With the visual feature pixel shuffling strategy, each image is further processed into a sequence of $169$ visual features. We limit the maximum number of images ($M$) in each sample to $50$, which produces up to $8,450$ visual features in total.
Models General Vision-Language Tasks Text-Rich Single
MIRB MiBench MMMU MathVista ScienceQA TextVQA DocVQA
Otter-9B 20.74 43.72 30.89 22.00 60.44 23.18 3.53
Emu2-Chat 36.02 58.93 34.10 30.40 65.69 66.60 5.44
VILA-LLaMA3-8B 40.87 53.70 36.90 35.40 79.90 66.30 30.38
mPlug-DocOwl-1.5 25.39 40.80 35.44 29.50 64.40 68.60 82.20
Idefics2-8B 33.02 46.39 42.90 45.00 89.04 70.40 67.30
LLaVA-NeXT-Inter 44.38 74.52 38.44 32.10 72.63 62.76 75.70
Llama-3.2-11B 20.96 34.33 50.70 51.50 71.99 73.60 88.40
Qwen2-VL-7B 59.96 52.91 54.10 58.20 78.38 84.30 94.50
Mantis-LLaVA 40.76 59.96 40.10 34.40 74.90 59.20 39.02
Mantis-Idefics2 41.80 56.80 41.10 40.40 81.30 63.50 54.03
Leopard-LLaVA 42.00 60.80 43.00 45.50 85.57 67.70 68.07
Leopard-Idefics2 41.38 61.74 40.11 44.80 90.38 80.40 74.79
Leopard-LLaVA-Pro 44.20 63.25 45.00 54.24 90.50 75.50 82.61
Table 7: Experimental results on out-of-domain evaluations, including general domain vision language tasks and text-rich single-image tasks.
Following Liu et al. (2023a), we adopt a two-layer MLPs as the visual-language connector. We use LLaMA-3.1 (Meta et al., 2024) as the backbone language model.
For Leopard-Idefics2, we follow the architecture of Idefics2-8B which uses SigLIP-SO-400M as the visual encoder but increases its image resolution to $980\times 980$ to make the text legible. The features outputted by the visual encoder are compressed with a feature resampler into $64$ tokens per image. Idefics2-8B adopts the Mistral-7B (Jiang et al., 2023) as the LM.
Training Details. When training Leopard-LLaVA, we first train the visual-language connector using LLaVA’s 558K multimodal pre-training dataset. Subsequently, we fine-tune the model (with both the connector and the LM unfrozen) using our Leopard-instruct data. As for Leopard-Idefics2, we directly finetune the pre-trained Idefics2 checkpoint on the Leopard-instruct dataset. To excel in text-rich scenarios, we additionally incorporate text-rich image data into the pre-training stage of Leopard-LLaVA, leading to the most functional model: Leopard-LLaVA-Pro. This includes Recap-COCO-30K (Li et al., 2024f), CC-3M (Changpinyo et al., 2021), Donut (Kim et al., 2022), the Cauldron (Laurençon et al., 2024b), and 4M Arxiv pages processed by an OCR toolkit [^3] https://github.com/facebookresearch/nougat.. We provide further training details in A.2.
@@ 4.2 Baseline Models
We compare Leopard against a wide range of open-source MLLMs that support multi-image inputs. These baseline models include Otter-9B (Li et al., 2023), Emu2-Chat-34B (Sun et al., 2023), Mantis (Jiang et al., 2024), VILA (Lin et al., 2023), Idefics2-8B (Laurençon et al., 2024b), LLaVA-NeXT-Interleave (Li et al., 2024c), Qwen2-VL (Wang et al., 2024), and LLaMA3.2 (Meta et al., 2024). Models that only support a single image input are excluded from our comparisons, except for mPlug-DocOwl-1.5 (Hu et al., 2024a), as it is primarily trained on visual document data and demonstrates strong capabilities on text-rich image tasks. Table 2 demonstrates a detailed comparison of the model training details of between baseline models and our proposed Leopard, which highlights their architecture, image resolution and training data differences.
@@ 4.3 Evaluating Benchmarks
We evaluated Leopard and baseline methods across three categories of vision-language tasks on (1) single text-rich image evaluation, (2) multiple text-rich images evaluation, and (3) general reasoning evaluation. Benchmarks for (1) include TextVQA (Singh et al., 2019b) and DocVQA (Mathew et al., 2021). Benchmarks for (2) include Multi-page DocVQA (Tito et al., 2022), DUDE (Landeghem et al., 2023), SlideVQA (Tanaka et al., 2023), Multihiertt (Zhao et al., 2022), and MultiChartQA (Zhu et al., 2024), which cover a diverse range of typical multi-image tasks, such as document understanding and slide question answering. Benchmarks for (3) include MMMU (Yue et al., 2024), MathVista (Lu et al., 2024), ScienceQA (Saikh et al., 2022), MIRB (Zhao et al., 2024) and MiBench (Liu et al., 2024b), which evaluate MLLMs from different perspectives, including world knowledge, mathematics, etc.
@@ 4.4 Main Experimental Results
Leopard achieves outstanding performance on text-rich, multi-image benchmarks, as shown in Table 4. Notably, both Leopard-LLaVA-Pro and Leopard-Idefics2 significantly outperform all baselines. Leopard-LLaVA-Pro becomes the strongest open-source MLLM in this area, achieving an average improvement of 8.46 points over the previous best.
In single-image text-rich scenarios shown in Table 7, Leopard outperforms several recent strong models, including VILA and LLaVA-NeXT. Leopard even achieves slightly higher average scores than the state-of-the-art mPlug model, despite mPlug being trained on 4M single-image data while Leopard is tuned on <200K. This demonstrates that training on multi-image data from Leopard-instruct also benefits model performance on single-image tasks.
@@ 4.5 More Analysis
Leopard-instruct mainly cover three main domains, i.e., documents & slides ( doc), tables & charts ( chart), and websites ( web). To assess the impact of data from different domains, we conduct ablation studies on three variants of Leopard, with the results presented in Table 6. Removing any part of the training data results in performance degradation. The most significant drop occurs when we exclude document data while removing web data leads to a slight decrease. Furthermore, eliminating all newly collected data (retaining only the existing 340K data) causes a substantial performance decline.
@@ 5 Conclusion
In this paper, we introduce Leopard, a novel MLLM specifically designed for text-rich, multi-image tasks. Leopard has two key innovations: (1) Leopard-instruct, a large-scale instruction-tuning dataset that encompasses a wide range of text-rich, multi-image instructions, and (2) an adaptive image encoding module capable of processing multiple high-resolution images efficiently. Our experimental results across diverse benchmarks highlight Leopard’s superior performance compared to existing open-source MLLMs, particularly in text-rich multi-image scenarios. Further analysis and ablation studies underscore the effectiveness of both the collected dataset and adaptive encoding strategy, solidifying Leopard’s contribution to multimodal research.
@@ A.1 Leopard-Instruct
To train Leopard, we created a large instruction-tuning dataset, Leopard-instruct, with 925K instances, including 739K designed for text-rich, multi-image scenarios. Despite surveying existing datasets, we found only 154K suitable text-rich, multi-image samples – insufficient for effective instruction tuning, which is far from sufficient for effective instruction tuning, as shown in prior MLLM studies (Jiang et al., 2024; Laurençon et al., 2024b; Li et al., 2024c). To overcome this limitation, we developed several data collection pipelines to collect high-quality text-rich, multi-image data, resulting in additional 585K instances.
@@
ArxivQA (Li et al., 2024d) Doc Reformed 1-3 Existing 81
DUDE (Landeghem et al., 2023) Doc Public 1-50 Augmented 23
MP-DocVQA (Tito et al., 2022) Doc Public 1-20 Augmented 36
DocVQA (Mathew et al., 2021) Doc No 1 None 39
TAT-DQA (Zhu et al., 2022) Doc Reformed 2-5 Augmented 13
SlidesGeneration (Sefid et al., 2021) Slides Repurposed 1-20 Augmented 3
SlidesVQA (Tanaka et al., 2023) Slides Public 20 Augmented 10
@@
Multihiertt (Zhao et al., 2022) Table Public 3-7 Existing/Augmented 15
MultiTabQA (Pal et al., 2023) Table Public 1-2 Augmented 6
TableGPT (Li et al., 2024e) Table Split 2 Existing 4
TabMWP (Lu et al., 2023) Table No 1 Existing 23
ChartGemma (Masry et al., 2024) Chart Reformed 1-4 Existing 65
DVQA (Kafle et al., 2018) Chart Reformed 1-3 None 200
FigureQA (Kahou et al., 2018) Chart Reformed 1-2 None 36
ChartQA (Masry et al., 2022) Chart Reformed 2 Augmented 32
@@
Mind2Web (Deng et al., 2023) Web Split 1-5 None 7
WebsiteScreenshots (Aydos, 2020) Web No 1 Augmented 2
Omniact (Kapoor et al., 2024) Web No 1 None 1
RICO (Hsiao et al., 2024) Web Reformed 1-4 None 25
WebVision (Li et al., 2017) Web No 1 Existing 1
WebUI (Wu et al., 2023a) Web No 1 None 19
LLaVAR (Zhang et al., 2023) Mix No 1 Existing 15
MathV360k (Shi et al., 2024) Mix No 1 None 38
Monkey (Li et al., 2024g) Mix Reformed 1-3 None 92
MPlugDocReason (Hu et al., 2024a) Mix No 1 Existing 25
IconQA (Lu et al., 2021) Other Public 1-6 Augmented 64
InfographicVQA (Mathew et al., 2022) Other No 1 Augmented 23
MapQA (Chang et al., 2022) Other Reformed 1-2 None 4
@@
A.2 Training Details
For Leopard-Idefics2, we note that the Idefics2 model is pre-trained on a dataset comprised of over 350M multimodal samples. Given the computational challenges of reproducing such extensive pre-training, and to ensure a fair comparison with baselines that utilize the pre-trained Idefics2 checkpoint, we directly adopt Idefics2’ visual feature resampler and fine-tune the model on the Leopard-instruct dataset.
We train both Leopard-LLaVA and Leopard-Idefics2 on 64 A100-40G GPUs with a global batch size of 128. We use the AdamW optimizer with $\beta_{1}=0.9$, $\beta_{2}=0.999$. Following (Jiang et al., 2024), we use a learning rate of $1\times 10^{-5}$ for Leopard-LLaVA and $5\times 10^{-6}$ for Leopard-Idefics2 to protect its pretrian knowledge. We use a cosine learning rate scheduler with a linear learning rate warm-up for the first $3\%$ steps. All model variants are trained 1 epoch under the same hyperparameters. It takes around 120 GPU days to train Leopard under both settings.
A.3 Prompts